OpenStudy (anonymous):

If (-3,0) is a point on the graph of y=f(x) which of the following must be on the graph of y=-f(x)? 1.)(-3,o) 2.)(3,0) 3.)(0,3) 4.)(3,3) 5.)(0,-3)

OpenStudy (anonymous):

I know that y=f(x) is the original point, and that y=-f(x) is its mirror point...and if the f is negate that means that its a reflection across the x axis...soooo...

OpenStudy (anonymous):

but theres no -3,3

OpenStudy (anonymous):

wouldn't it be 1

OpenStudy (anonymous):

but thats the original point

OpenStudy (anonymous):

cann these work like a regular equation? like y=mx+b cuz if so you don't do anything to the y-intercept you just change the slope it will go through the same point. or no? i just haven't done these with functions

OpenStudy (anonymous):

im not sure, not much of that makes sense and my book hasn't talked about it.

OpenStudy (anonymous):

the lesson was just on reflections across the point of origin, x axis, and y axis.

OpenStudy (anonymous):

Well if that is true. then it should be 1.

OpenStudy (anonymous):

answer choice 1 that is

OpenStudy (anonymous):

How can the original point of (-3,0) reflect across the x axis and still be the point (-3,0)? thats no clicking my brain

OpenStudy (anonymous):



OpenStudy (anonymous):

now i've gotta go school lol. later.

OpenStudy (anonymous):

Where is (-3,0)? - The 0 is a clue!

OpenStudy (anonymous):